# Stage-wise Dynamics of Classifier-Free Guidance in Diffusion Models

### v1

@@ 1 Introduction
Diffusion models have driven remarkable progress in generative modeling, achieving state-of-the-art results across images, video, audio, and multimodal domains (Rombach et al., 2022; Podell et al., 2023; Cao et al., 2024; Zhang et al., 2025). A central requirement in practice is conditional generation, where outputs must follow prompts, class labels, or other structured signals. Among existing strategies, Classifier-Free Guidance (CFG) (Ho & Salimans, 2021) has become the de facto standard due to its simplicity and effectiveness, powering nearly all large-scale diffusion pipelines. Numerous variants further extend its influence (Jin et al., 2025; Gao et al., 2025; Malarz et al., 2025; Chung et al., 2025; Sadat et al., 2025; Castillo et al., 2025). Yet despite its widespread adoption, the theoretical underpinnings of CFG remain poorly understood.
While recent work has advanced the theory of diffusion sampling (Benton et al., 2023; Chen et al., 2023; Cai & Li, 2025; Li et al., 2025a), the analysis of CFG itself is still in its infancy, hindered by its heuristic formulation. Prior studies fall into two categories. The first assumes a unimodal conditional distribution—typically a single Gaussian—yielding clean derivations but overlooking the multimodal nature of real-world tasks (Chidambaram et al., 2024; Wu et al., 2024; Bradley & Nakkiran, 2024; Xia et al., 2024; Jin et al., 2025; Pavasovic et al., 2025; Li et al., 2025b). The second relaxes these assumptions but imposes only weak conditions, producing broad qualitative insights without sharp predictions (Li & Jiao, 2025). Consequently, several well-documented empirical phenomena remain theoretically elusive, most notably the collapse of diversity under large guidance weights (Ho & Salimans, 2021). Explaining this diversity loss is both a key theoretical challenge and a practical necessity for improving conditional generation.
@@
We empirically validate these predictions on state-of-the-art diffusion models rather than toy settings. To test our framework, we first compare early-high, late-high, and all-high schedules, confirming that strong early guidance reduces global diversity. We then probe the late stage by initializing trajectories with identical noise and perturbations, finding that large late weights enforce excessive similarity in fine details and diminish local diversity. Beyond these validation studies, we further conduct an additional experiment with a time-varying guidance schedule. The results show that adjusting guidance strength over time can yield improved overall outcomes—a practical insight that, while secondary to our main theoretical contribution, highlights the broader value of our analysis. Our implementation is publicly available at https://github.com/sqt24/tvcfg.
@@ 2.1 Diffusion Model
where $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ is the score function, representing the gradient of the log-density, and $\bar{{\bm{w}}}_{t}$ is a backward Wiener process. Intuitively, the score term provides a restoring drift that counteracts noise, gradually steering trajectories back toward high-density regions of the data distribution.
Modern implementations typically adopt the equivalent probability flow ODE (Song et al., 2021), which eliminates randomness while preserving the same marginal distributions:
@@ 3 Stage-wise Behavior of Classifier Free Guidance in Sampling
(A1) Unconditional Distribution. $p_{0}({\bm{x}})=\mathcal{N}(\mathbf{0},{\bm{I}}_{d})$. This aligns with the standard Gaussian prior used in latent diffusion models, making it both realistic and analytically convenient.
@@ 3.1 First Stage: Acceleration and Direction Shift
Let $\bar{\bm{\mu}}=\sum_{k=1}^{K}\pi_{k}{\bm{\mu}}_{k}$ denote the class-weighted mean of the Gaussian mixture prior, and assume the same initialization ${\bm{x}}_{1}\sim\mathcal{N}(\mathbf{0},{\bm{I}})$ for both trajectories. Then for any $\omega>1$, there exists a time point $t_{e1}<1$ such that, for all $t\in[t_{e1},1)$,
@@
where ${\bm{x}}_{t}^{(y)}$ is the solution to the conditional probability flow ODE and ${\bm{x}}_{t}^{(\textsc{CFG})}$ the solution to the CFG-driven ODE, both starting from ${\bm{x}}_{1}$. In expectation, the CFG-driven trajectory remains closer to the reference point $\omega\bar{\bm{\mu}}$ than the purely conditional one in the early stage of sampling.
@@ 3.2 Second Stage: Intra-class Mode Separation
$p_{0}({\bm{x}}\mid y)=\pi_{1}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{1},\sigma^{2}{\bm{I}})+\pi_{2}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{2},\sigma^{2}{\bm{I}}),\quad\|{\bm{\mu}}_{1}\|=\|{\bm{\mu}}_{2}\|,\;\pi_{1}<\pi_{2}.$
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big)>\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big)\quad\forall\,t\in[0,t_{s_{2}}].$
@@
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big)<\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\bm{I}}\big).$
@@
Proposition 3.4 makes explicit how the first stage biases the sampling distribution. Once a trajectory is drawn sufficiently close to the scaled mean $k\bar{\bm{\mu}}$, the surrounding drift field guarantees that the posterior likelihood of ${\bm{\mu}}_{2}$ will dominate that of ${\bm{\mu}}_{1}$ at all earlier times. As $k$ increases with $\omega$, the neighborhood where this dominance holds also expands. Consequently, although the basin of ${\bm{\mu}}_{1}$ mathematically persists, it becomes increasingly difficult for trajectories to reach it under strong guidance. The scarcity of weaker-mode samples in practice therefore reflects not the elimination of $U_{s_{2}}$, but the fact that most trajectories are already preconditioned by the first stage to fall outside it.
@@ 3.3 Third Stage: Concentration
Under mild assumption, there exist a time $t_{s_{3}}\in(0,1)$ and a radius $r>0$ such that the following holds uniformly. For any $k\in\{1,2\}$ and any initial pair
@@ 3.4 Summary across stages
Taken together, these stages explain the dual empirical effects of CFG: it improves conditional fidelity and visual sharpness, yet simultaneously diminishes both global coverage (loss of weaker modes) and local diversity (within-mode contraction). This stage-wise perspective provides a unified theoretical framework for understanding diversity loss under CFG and points to principled strategies for designing time-varying guidance schedules.
@@ 4 Experiments
(a) Constant high weight
(b) Early low weight
(c) Early high weight
Figure 2: Comparison of guidance schedules on the prompt “A view of a bathroom that is clean.”. (a) Constant schedule and (c) Early-high schedule both collapse diversity, with most samples converging to layouts dominated by large windows and uniform cool tones. (b) Early-low schedule mitigates this effect, producing more varied spatial structures and color palettes.
@@
(a) Late-high schedule, MSE=0.0025
(b) Constant schedule, MSE=0.0103
4.2 Methodological implications.
Guided by the above analysis, we propose a time-varying schedule where the guidance weight rises early, peaks at the intermediate stage, and decreases late. Unlike heuristic on–off schemes (Kynkäänniemi et al., 2024), it evolves smoothly without abrupt switches. Though not our main focus, this theory-derived design yields empirical gains and highlights how stage-wise insights translate into practical improvements in robustness and diversity.
@@
(d) vanilla-APG
(e) interval-APG
(f) TV-APG
Figure 4: Generated samples with the prompt “A view of a bathroom that is clean” at high sampling budget (NFE=50). Constant schedules yield semantically consistent but overly uniform results, while interval schedules improve diversity at the cost of semantic stability (e.g., red stain in the 2nd row, 4th column; monotonic patterns in the 3rd row, 4th column). By contrast, TV schedules achieve both higher diversity and stronger semantic alignment.
Figure 4 shows that our schedule preserves intra-class diversity while reducing over-saturation. Table 1 further reports consistent gains in ImageReward (IR) (Xu et al., 2024), which better reflects overall quality than CLIP. At low weights our method matches vanilla CFG, but at high weights—where vanilla CFG degrades—it achieves clear advantages and stronger best-case performance. The same principle applies to APG (Sadat et al., 2025), where TV-APG likewise improves IR.
@@
Table 1: Results with NFE fixed at 1010 across guidance scales. Bold indicates the best value at each $\omega$ across variants. The Best column reports, for each variant, its optimum over $\omega$.
Metric Variant CFG APG $\omega{=}3$ $\omega{=}5$ $\omega{=}7$ $\omega{=}9$ Best $\omega{=}3$ $\omega{=}5$ $\omega{=}7$ $\omega{=}9$ Best IR $\uparrow$ vanilla 0.894 0.806 0.553 0.223 0.894 0.909 0.890 0.712 0.434 0.909 interval 0.602 0.694 0.727 0.723 0.727 0.602 0.693 0.724 0.725 0.725 TV(Ours) 0.859 0.935 0.950 0.932 0.950 0.860 0.937 0.958 0.941 0.958 CLIP $\uparrow$ vanilla 0.320 0.322 0.319 0.313 0.322 0.322 0.322 0.321 0.317 0.322 interval 0.315 0.317 0.318 0.319 0.319 0.315 0.317 0.318 0.319 0.319 TV(Ours) 0.319 0.319 0.319 0.319 0.319 0.319 0.320 0.320 0.319 0.320 FID $\downarrow$ vanilla 28.305 29.275 32.859 38.988 28.305 28.028 28.208 30.268 34.950 28.028 interval 30.485 28.167 27.884 28.413 27.884 30.461 28.155 28.000 28.427 28.000 TV(Ours) 27.898 27.722 28.547 30.259 27.722 27.917 27.935 28.503 29.838 27.917 Saturation vanilla 0.283 0.408 0.513 0.574 N/A 0.266 0.385 0.500 0.576 N/A interval 0.173 0.178 0.185 0.193 N/A 0.172 0.177 0.183 0.192 N/A TV(Ours) 0.197 0.229 0.263 0.298 N/A 0.193 0.221 0.253 0.288 N/A
Table 2: Results with guidance weight fixed at $\omega=9$ while varying the NFE budget. Bold indicates the best value within each method at the same NFE.
Metric Variant CFG APG $\text{NFE}=5$ $\text{NFE}=10$ $\text{NFE}=15$ $\text{NFE}=20$ $\text{NFE}=5$ $\text{NFE}=10$ $\text{NFE}=15$ $\text{NFE}=20$ IR $\uparrow$ vanilla -1.137 0.223 0.616 0.820 -0.926 0.434 0.735 0.860 interval 0.145 0.723 0.827 0.863 0.145 0.725 0.825 0.865 TV(Ours) 0.176 0.932 1.016 1.049 0.174 0.941 1.029 1.061 CLIP $\uparrow$ vanilla 0.275 0.313 0.319 0.320 0.282 0.317 0.320 0.320 interval 0.314 0.319 0.318 0.317 0.314 0.319 0.319 0.317 TV(Ours) 0.312 0.319 0.319 0.319 0.312 0.319 0.319 0.319 FID $\downarrow$ vanilla 80.482 38.988 31.863 29.059 73.751 34.950 30.122 28.602 interval 47.239 28.413 25.747 25.385 47.735 28.427 25.691 25.365 TV(Ours) 46.895 30.259 29.256 28.458 46.878 29.838 28.834 27.965 Saturation vanilla 0.589 0.574 0.526 0.476 0.610 0.576 0.521 0.472 interval 0.197 0.193 0.212 0.233 0.194 0.192 0.209 0.229 TV(Ours) 0.271 0.298 0.325 0.332 0.268 0.288 0.309 0.311

### v2

@@ 1 Introduction
Diffusion models have driven remarkable progress in generative modeling, achieving state-of-the-art results across images, video, audio, and multimodal domains (Rombach et al., 2022; Podell et al., 2023; Cao et al., 2024; Zhang et al., 2025). A central requirement in practice is conditional generation, where outputs must follow prompts, class labels, or other structured signals. Among existing strategies, Classifier-Free Guidance (CFG) (Ho and Salimans, 2021) has become the de facto standard due to its simplicity and effectiveness, powering nearly all large-scale diffusion pipelines. Numerous variants further extend its influence (Jin et al., 2025; Gao et al., 2025; Malarz et al., 2025; Chung et al., 2025; Sadat et al., 2025; Castillo et al., 2025). Yet despite its widespread adoption, the theoretical underpinnings of CFG remain poorly understood.
While recent work has advanced the theory of diffusion sampling (Benton et al., 2023; Chen et al., 2023; Cai and Li, 2025; Li et al., 2025a), the analysis of CFG itself is still in its infancy, hindered by its heuristic formulation. Prior studies fall into two categories. The first assumes a unimodal conditional distribution—typically a single Gaussian—yielding clean derivations but overlooking the multimodal nature of real-world tasks (Chidambaram et al., 2024; Wu et al., 2024; Bradley and Nakkiran, 2024; Xia et al., 2024; Jin et al., 2025; Pavasovic et al., 2025; Li et al., 2025b). The second relaxes these assumptions but imposes only weak conditions, producing broad qualitative insights without sharp predictions (Li and Jiao, 2025). Consequently, several well-documented empirical phenomena remain theoretically elusive, most notably the collapse of diversity under large guidance weights (Ho and Salimans, 2021). Explaining this diversity loss is both a key theoretical challenge and a practical necessity for improving conditional generation. A detailed discussion of these related works can be found in Appendix.
@@
We empirically validate these predictions on state-of-the-art diffusion models rather than toy settings. To test our framework, we first compare early-high, late-high, and all-high schedules, confirming that strong early guidance reduces global diversity. To probe the late-stage behavior, we initialize trajectories from identical noise and introduce small perturbations during the intermediate stage. We observe that excessively large late-stage guidance forces trajectories to converge toward nearly identical fine details, thereby diminishing local diversity. Beyond these validation studies, we further conduct an additional experiment with a time-varying guidance schedule which can improve performance. Although this is not the central theoretical contribution of this work, it underscores the practical relevance of our analysis. Our implementation is publicly available at https://github.com/sqt24/tvcfg.
@@ 2.1 Diffusion Model
where $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ is the score function, representing the gradient of the log-density, and $\bar{{\bm{w}}}_{t}$ is a backward Wiener process.
Modern implementations typically adopt the probability flow ODE (Song et al., 2021), which eliminates randomness while preserving the same marginal distributions:
@@ 3 Stage-wise Behavior of Classifier Free Guidance in Sampling
(A1) Unconditional Distribution. $p_{0}({\bm{x}})=\mathcal{N}(\mathbf{0},{\mathbf{I}}_{d})$. This aligns with the standard Gaussian prior used in latent diffusion models, making it both realistic and analytically convenient.
@@ 3.1 First Stage: Acceleration and Direction Shift
Let $\bar{\bm{\mu}}=\sum_{k=1}^{K}\pi_{k}{\bm{\mu}}_{k}$ denote the class-weighted mean of the Gaussian mixture prior, and assume the same initialization ${\bm{x}}_{1}\sim\mathcal{N}(\mathbf{0},{\mathbf{I}})$ for both trajectories. Then for any $\omega>1$, there exists a time point $t_{e1}<1$ such that, for all $t\in[t_{e1},1)$,
@@
where $\omega$ is the guidance weight, ${\bm{x}}_{t}^{(y)}$ is the solution to the conditional probability flow ODE and ${\bm{x}}_{t}^{(\textsc{CFG})}$ the solution to the CFG-driven ODE, both starting from ${\bm{x}}_{1}$. In expectation, the CFG-driven trajectory remains closer to the reference point $\omega\bar{\bm{\mu}}$ than the purely conditional one in the early stage of sampling.
@@ 3.2 Second Stage: Intra-class Mode Separation
$p_{0}({\bm{x}}\mid y)=\pi_{1}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{1},\sigma^{2}{\mathbf{I}})+\pi_{2}\,\mathcal{N}({\bm{x}};{\bm{\mu}}_{2},\sigma^{2}{\mathbf{I}}),\quad\|{\bm{\mu}}_{1}\|=\|{\bm{\mu}}_{2}\|,\;\pi_{1}<\pi_{2}.$
Under some mild assumptions, there exist $t_{s_{2}}\in(0,1)$ and an $\omega$-independent region $U_{s_{2}}\subset\mathbb{R}^{d}$, depending only on $(\pi_{k},{\bm{\mu}}_{k},\sigma)$, such that for any $\omega\geq 1$, if ${\bm{x}}_{t_{s_{2}}}\in U_{s_{2}}$ and ${\bm{x}}_{t}$ follows the CFG probability flow ODE equation 6 on $0,t_{s_{2}}]$, then
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big)>\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big)\quad\forall\,t\in[0,t_{s_{2}}].$
@@
$\pi_{1}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{1},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big)<\pi_{2}\,\mathcal{N}\!\big({\bm{x}}_{t};(1-t){\bm{\mu}}_{2},\,(t^{2}+(1-t)^{2}\sigma^{2}){\mathbf{I}}\big).$
@@
Proposition 3.4 makes explicit how the first stage biases the sampling distribution. Once a trajectory is drawn sufficiently close to the scaled mean $k\bar{\bm{\mu}}$, the surrounding drift field guarantees that the posterior likelihood of ${\bm{\mu}}_{2}$ will dominate that of ${\bm{\mu}}_{1}$ at all earlier times. According to Theorem 3.2, a larger $\omega$ causes early-stage trajectories to be closer to $\omega\bar{{\bm{\mu}}}$, corresponding to a larger effective $k$ and thus expanding the region dominated by the stronger mode. Consequently, although the basin of ${\bm{\mu}}_{1}$ mathematically persists, it becomes increasingly difficult for trajectories to reach it under strong guidance. The scarcity of weaker-mode samples in practice therefore reflects not the elimination of $U_{s_{2}}$, but the fact that most trajectories are already preconditioned by the first stage to fall outside it.
@@ 3.3 Third Stage: Concentration
Under some mild assumptions, there exist a time $t_{s_{3}}\in(0,1)$ and a radius $r>0$ such that the following holds uniformly. For any $k\in\{1,2\}$ and any initial pair
@@
(a) Constant high weight
(b) Early low weight
(c) Early high weight
Figure 2: Comparison of guidance schedules on the prompt “A view of a bathroom that is clean.” The (a) Constant schedule and (c) Early-high schedule both collapse diversity, with most samples converging to layouts dominated by large windows and uniform cool tones. The (b) Early-low schedule mitigates this effect, producing more varied spatial structures and color palettes.
@@ 3.4 Summary across stages
Taken together, these stages explain the dual empirical effects of CFG: it improves conditional fidelity and visual sharpness, yet simultaneously diminishes both global coverage (loss of weaker modes) and local diversity (within-mode contraction). Interestingly, this resonates with prior work (Balaji et al., 2022; Li et al., 2024) that empirically found diffusion models determine the overall shape early in the process and fine details late, which corresponds to our distinction between global diversity and local diversity. This stage-wise perspective provides a unified theoretical framework for understanding diversity loss under CFG and points to principled strategies for designing time-varying guidance schedules.
(a) Late-high schedule, MSE=0.0025
(b) Constant schedule, MSE=0.0103
@@ 4 Experiments
Table 1: Results with NFE fixed at 1010 across guidance scales. Bold indicates the best value within each method at the same $\omega$ (higher is better for CLIP/IR; lower is better for FID).
Metric Variant CFG APG $\omega{=}3$ $\omega{=}5$ $\omega{=}7$ $\omega{=}9$ $\omega{=}3$ $\omega{=}5$ $\omega{=}7$ $\omega{=}9$ CLIP $\uparrow$ vanilla 0.320 0.322 0.319 0.313 0.322 0.322 0.321 0.317 interval 0.315 0.317 0.318 0.319 0.315 0.317 0.318 0.319 TV(Ours) 0.319 0.319 0.319 0.319 0.319 0.320 0.320 0.319 $\beta$ 0.316 0.318 0.318 0.316 0.317 0.318 0.318 0.317 IR $\uparrow$ vanilla 0.894 0.806 0.553 0.223 0.909 0.890 0.712 0.434 interval 0.602 0.694 0.727 0.723 0.602 0.693 0.724 0.725 TV(Ours) 0.859 0.935 0.950 0.932 0.860 0.937 0.958 0.941 $\beta$ 0.705 0.822 0.831 0.790 0.737 0.846 0.867 0.833 FID $\downarrow$ vanilla 28.305 29.275 32.859 38.988 28.028 28.208 30.268 34.950 interval 30.485 28.167 27.884 28.413 30.461 28.155 28.000 28.427 TV(Ours) 27.898 27.722 28.547 30.259 27.917 27.935 28.503 29.838 $\beta$ 28.679 28.571 29.774 32.184 28.163 28.082 29.148 31.615 Saturation vanilla 0.283 0.408 0.513 0.574 0.266 0.385 0.500 0.576 interval 0.173 0.178 0.185 0.193 0.172 0.177 0.183 0.192 TV(Ours) 0.197 0.229 0.263 0.298 0.193 0.221 0.253 0.288 $\beta$ 0.184 0.219 0.259 0.305 0.186 0.218 0.257 0.304 Diversity $\downarrow$ vanilla 1.066 1.101 1.105 1.081 1.059 1.115 1.136 1.128 interval 1.013 1.073 1.122 1.160 1.013 1.072 1.121 1.159 TV(Ours) 1.092 1.158 1.196 1.223 1.088 1.154 1.194 1.222 $\beta$ 1.123 1.181 1.217 1.242 1.121 1.178 1.214 1.241
@@
(d) $\beta$-CFG
Figure 4: Generated samples with the prompt “A view of a bathroom that is clean” at high sampling budget (NFE=50). While constant schedules yield semantically consistent but overly uniform results, methods adhering to the low-high-low scheduling principle exhibit significantly higher diversity.
Table 2: Results with guidance weight fixed at $\omega=9$ while varying the NFE budget. Bold indicates the best value within each method at the same NFE (higher is better for CLIP/IR; lower is better for FID).
Metric Variant CFG APG $\text{NFE}=5$ $\text{NFE}=10$ $\text{NFE}=15$ $\text{NFE}=20$ $\text{NFE}=5$ $\text{NFE}=10$ $\text{NFE}=15$ $\text{NFE}=20$ CLIP $\uparrow$ vanilla 0.275 0.313 0.319 0.320 0.282 0.317 0.320 0.320 interval 0.314 0.319 0.318 0.317 0.314 0.319 0.319 0.317 TV(Ours) 0.312 0.319 0.319 0.319 0.312 0.319 0.319 0.319 $\beta$ 0.302 0.316 0.318 0.318 0.304 0.317 0.318 0.318 IR $\uparrow$ vanilla -1.137 0.223 0.616 0.820 -0.926 0.434 0.735 0.860 interval 0.145 0.723 0.827 0.863 0.145 0.725 0.825 0.865 TV(Ours) 0.176 0.932 1.016 1.049 0.174 0.941 1.029 1.061 $\beta$ -0.328 0.790 0.971 1.024 -0.202 0.833 0.984 1.036 FID $\downarrow$ vanilla 80.482 38.988 31.863 29.059 73.751 34.950 30.122 28.602 interval 47.239 28.413 25.747 25.385 47.735 28.427 25.691 25.365 TV(Ours) 46.895 30.259 29.256 28.458 46.878 29.838 28.834 27.965 $\beta$ 62.979 32.184 29.246 28.348 58.398 31.615 28.836 27.968 Saturation vanilla 0.589 0.574 0.526 0.476 0.610 0.576 0.521 0.472 interval 0.197 0.193 0.212 0.233 0.194 0.192 0.209 0.229 TV(Ours) 0.271 0.298 0.325 0.332 0.268 0.288 0.309 0.311 $\beta$ 0.236 0.305 0.316 0.330 0.243 0.304 0.307 0.315 Diversity $\uparrow$ vanilla 0.867 1.081 1.171 1.201 0.917 1.128 1.186 1.200 interval 1.140 1.160 1.178 1.202 1.140 1.159 1.176 1.200 TV(Ours) 1.191 1.223 1.233 1.231 1.191 1.222 1.232 1.231 $\beta$ 1.196 1.242 1.241 1.239 1.198 1.241 1.239 1.238
4.2 Methodological implications.
Guided by the above analysis, we propose a linear time-varying schedule where the guidance weight rises early, peaks at the intermediate stage, and decreases late (TV-CFG). We also considered Interval-CFG (Kynkäänniemi et al., 2024) and $\beta$-CFG methods (Malarz et al., 2025). Both adhere to the time-varying scheduling principle of being weak in the early stages, strong in the middle, and weak in the late stages, which theoretically contributes to sampling diversity. The specific scheduling principles are detailed in the Appendix. Though not our main focus, this theory-derived design yields empirical gains and highlights how stage-wise insights translate into practical improvements in robustness and diversity.
Figure 4 shows that our schedule preserves diversity while reducing over-saturation. Table 1 further reports consistent gains in ImageReward (IR) (Xu et al., 2024), which better reflects overall quality than CLIP. Notably, all three methods that follow our design principle effectively improve generation diversity while maintaining good generation quality. At low weights our method matches vanilla CFG, but at high weights—where vanilla CFG degrades—it achieves clear advantages and stronger best-case performance. The same principle applies to APG (Sadat et al., 2025), where TV-APG likewise improves IR.